# Control of a commercial vehicle by a tetraplegic human using a bimanual brain-computer interface

### v1

@@ Acknowledgments
Control of a commercial vehicle by a tetraplegic human using a bimanual brain-computer interface
@@
Figure S1: Brain implant locations of BCI-$J$. We implanted $1\times 96$ channels of NeuroPort Electrodes (Blackrock Neurotech, Salt Lake City, UT) in his posterior parietal cortex (PPC) and $1\times 96$ channels of NeuroPort Electrodes in the hand knob region of his motor cortex (MC). “CS” stands for the central sulcus. “IPS” stands for the intraparietal sulcus.
@@
Figure S2: Trial-based reaction accuracy for the simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We applied pairwise comparisons with the Bonferroni correction from a multiple comparison test using the information contained from one-way ANOVA among 21 participants. For participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of reaction accuracy among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
@@
Figure S3: Trial-based reaction sensitivity for the simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We applied pairwise comparisons with the Bonferroni correction from a multiple comparison test using the information contained from one-way ANOVA among 21 participants. For participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of reaction sensitivity among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
@@
Figure S4: Trial-based reaction specificity for the simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We applied pairwise comparisons with the Bonferroni correction from a multiple comparison test using the information contained from one-way ANOVA among 21 participants. For participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of reaction specificity among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
@@
Table S6: Two-sample $t$-test $p$ values for trial-based simple reaction metrics, when equal variances were not assumed, between BCI-JJ and the motor intact control group with the same right index finger effector. We collected 10 runs from BCI-JJ and 4 runs from each of the 20 motor intact participants whose average age was the same as the age of BCI-$J$. There was no gender bias within the motor intact control group. During each run, there were 40 GO trials and 10 randomly interleaved NO-GO catch trials. The valid reaction time (RT) ranged from 50 ms to 1000 ms in GO trials. We used a left-tailed $t$-test to test against the alternative hypothesis that the population mean of the valid RT for BCI-JJ was less than that for the motor intact control group of 20 participants. We used a both-tailed $t$-test to test against the alternative hypothesis that the population means for each performance measure were not equal between BCI-JJ and the motor intact control group. If $p<0.05$, the metrics had a significant impact on the performance.
@@
Table S7: Pairwise comparison $p$ values for trial-based simple reaction metrics between BCI-JJ and each of the 20 motor intact control participants with the same right index finger effector. The $p$ values were obtained from a multiple comparison test using the results of one-way ANOVA. We collected 10 runs from BCI-JJ and 4 runs from each motor intact participant (M01–M20). In the parenthesis after each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The valid reaction time (RT) ranged from 50 ms to 1000 ms in GO trials. Bonferroni correction for multiple comparisons was applied by setting the significance level to the corrected $p<0.05$.
@@
Table S11: Simulated driving scores and infractions, mean (SD), and two-sample $t$-test $p$ values between BCI-JJ and the motor intact control group using the same bimanual control. We collected 10 runs from BCI-JJ and 2 runs from each of the 20 motor intact participants whose average age was the same as the age of BCI-$J$. There was no gender bias within the motor intact control group. During each run of navigating a virtual vehicle in a simulated town environment with busy traffic and a route modified from the CARLA Leaderboard 2.0, a participant used the right thumb for cursor movement to control its steering and speed and the left index finger for clicks to control its full-stop braking. The simulated driving scores ranged from 0 to 1.0. The proportion of the completed route distance (i.e., $C$ for route completion) could be a decimal between 0 and 1.0 with 1.0 as full completion of the route. The rest infraction components, which described the occurrences of collisions, running red lights, and lane deviations (i.e., $N_{c}$, $N_{l}$ and $N_{s}$ respectively) per run, were all nonnegative numbers. We used a right-tailed $t$-test to test against the alternative hypothesis that the population mean of the simulated driving score for BCI-JJ was greater than that for the motor intact control group of 20 participants. We used a both-tailed $t$-test to test against the alternative hypothesis that the population means for each infraction component were not equal between BCI-JJ and the motor intact control group. If $p<0.05$, the metrics had a significant impact on the performance.
@@
Caption for Movie S1: BCI teledriving of a commercial Ford Mustang Mach-E via Route 1 in Mcity by BCI-JJ using the unilateral cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle (bottom-left) would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 1 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S2: BCI teledriving of a commercial Ford Mustang Mach-E via Route 2 in Mcity by BCI-JJ using the unilateral cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle (bottom-left) would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 2 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S3: BCI teledriving of a commercial Ford Mustang Mach-E via Route 3 in Mcity by BCI-JJ using the unilateral cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 3 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S4: BCI teledriving of a commercial Ford Mustang Mach-E via Route 4 in Mcity by BCI-JJ using the unilateral cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. The cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle (bottom-left) would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 4 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
Caption for Movie S5: Side-by-side comparisons of the average simulated town driving performance between BCI-JJ and a motor intact participant using the bimanual cursor-and-click control. We showed two example runs from BCI-JJ and one motor intact participant on the town route modified from the CARLA Leaderboard 2.0. Among all the collected runs, these two example runs had their simulated driving scores closest to the average scores for BCI-JJ and the motor intact control group. We retrieved the cursor movement and click information from either the BCI decoder for BCI-JJ or the joystick for motor intact participants. With the overlay instance floating above the vehicle simulation window, each participant got real-time visual feedback of their cursor and click control, the view ahead of the vehicle, and the side views. The cursor movement with the right thumb controlled steering and speed of the virtual vehicle, whereas the clicks with the left index finger controlled its full-stop braking. Each route map (top) corresponding to each participant’s driving recording has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S6: BCI teledriving of a commercial Ford Mustang Mach-E through the obstacle course clockwise with a lane-switch segment by BCI-JJ using the bimanual cursor-and-click control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at 2 times the speed of the original recording.
@@
Caption for Movie S7: BCI teledriving of a commercial Ford Mustang Mach-E through the obstacle course clockwise without any lane switch by BCI-JJ using the bimanual cursor-and-click control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at 2 times the speed of the original recording.
@@
Caption for Movie S8: BCI teledriving of a commercial Ford Mustang Mach-E through the obstacle course counterclockwise with a lane-switch segment by BCI-JJ using the bimanual cursor-and-click control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at 2 times the speed of the original recording.
@@
Caption for Movie S9: BCI teledriving of a commercial Ford Mustang Mach-E through the obstacle course counterclockwise without any lane switch by BCI-JJ using the bimanual cursor-and-click control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at 2 times the speed of the original recording.

### v2

@@ Acknowledgments
Control of a commercially available vehicle by a tetraplegic human using a brain-computer interface
@@
Figure S1: Brain implant locations of BCI-$J$. We implanted one array of 96-channel NeuroPort Electrodes © Blackrock Neurotech in the left superior parietal lobule of his posterior parietal cortex (PPC) (MNI -34, -58, 67) and the other array of 96-channel NeuroPort Electrodes in his left motor cortex (MC) close to the hand knob (MNI -39, -15, 69). CS: central sulcus. IPS: intraparietal sulcus.
@@
Figure S2: Trial-based valid reaction times for simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We conducted pairwise comparisons using a Bonferroni post hoc test following a one-way ANOVA among 21 participants. For the participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of the valid reaction times between 50 ms and 1000 ms in GO trials/phases among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
@@
Figure S3: Trial-based reaction accuracy for simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We conducted pairwise comparisons using a Bonferroni post hoc test following a one-way ANOVA among 21 participants. For participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of reaction accuracy among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
@@
Figure S4: Trial-based reaction sensitivity for simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We conducted pairwise comparisons using a Bonferroni post hoc test following a one-way ANOVA among 21 participants. For participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of reaction sensitivity among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
A
B
Figure S5: Trial-based reaction specificity for simple and braking reaction time tasks among BCI-JJ and all 20 motor intact participants individually with the same right index finger effector. In the parenthesis under each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The box plots are sorted according to the average performance of each participant. We conducted pairwise comparisons using a Bonferroni post hoc test following a one-way ANOVA among 21 participants. For participants whose box plots are outside the green rectangular area, their performance was different from BCI-JJ at the corrected $5\%$ significance level. There are comparisons of reaction specificity among all 21 participants for (A) the simple reaction time task and (B) the braking reaction time task.
@@
A
B
C
D
E
F
Figure S8: Results from a post-session questionnaire collected from 11 out of 20 motor intact participants after they completed the simulated town driving task. (A) Whether they usually experience motion sickness in a physical vehicle. (B) Whether they experienced motion sickness during this task. (C) Whether they had any past experience(s) in driving in a simulated environment before participating in this task. (D) Whether they had any past experience(s) in controlling a gaming joystick (similar to the one they used in this task) before participating in this task. (E) How much difficulty they felt to control the gaming joystick during this task. (F) Whether they were fatigued by this task.
A
B
C
Figure S9: Video lag analysis for three driving tasks. (A) Boxplot of the lags between intention and video feedback during BCI driving sessions grouped by task, including the Mcity teledriving task, the obstacle-course teledriving task, and the simulated town driving task. Intention was obtained from the overlay state, and execution was calculated from the movement of the vehicle in videos. (B) Segment of the originally extracted steering intention and video feedback values from video recording in the Mcity teledriving task. (C) Aligned steering intention and video feedback values for the segment in (B) with an estimated lag of -193 ms.
@@
Table S6: Two-sample $t$-test $p$ values for trial-based simple reaction metrics, when equal variances were not assumed, between BCI-JJ and the motor intact control group with the same right index finger effector. We collected 10 runs from BCI-JJ and 4 runs from each of the 20 motor intact participants whose average age was the same as the age of BCI-$J$. There was no gender bias within the motor intact control group. During each run, there were 40 GO trials and 10 randomly interleaved NO-GO catch trials. The valid reaction time (RT) ranged from 50 ms and 1000 ms in GO trials. We used a left-tailed $t$-test to test against the alternative hypothesis that the population mean of the valid RT for BCI-JJ was less than that for the motor intact control group of 20 participants. We used a both-tailed $t$-test to test against the alternative hypothesis that the population means for each performance measure were not equal between BCI-JJ and the motor intact control group. If $p<0.05$, the metrics had a significant impact on the performance.
@@
Table S7: Pairwise comparison $p$ values for trial-based simple reaction metrics between BCI-JJ and each of the 20 motor intact control participants with the same right index finger effector. The $p$ values were obtained from a multiple comparison test using the results of one-way ANOVA. We collected 10 runs from BCI-JJ and 4 runs from each motor intact participant (M01–M20). In the parenthesis after each participant’s ID, the number represents their age, and the letter “f” or “m” represents their gender (female or male). The valid reaction time (RT) ranged from 50 ms and 1000 ms in GO trials. Bonferroni correction for multiple comparisons was applied by setting the significance level to the corrected $p<0.05$.
@@
Table S11: Simulated driving scores and infractions, mean (SD), and two-sample $t$-test $p$ values between BCI-JJ and the motor intact control group using the same bimanual control. We collected 10 runs from BCI-JJ and 2 runs from each of the 20 motor intact participants whose average age was the same as the age of BCI-$J$. There was no gender bias within the motor intact control group. During each run of navigating a virtual vehicle in a simulated town environment with heavy traffic and a route modified from the CARLA Leaderboard 2.0, a participant used the right thumb for cursor movement to control its steering and speed and the left index finger for clicks to control its full-stop braking. The simulated driving scores ranged from 0 to 1.0. The proportion of the completed route distance (i.e., $C$ for route completion) could be a decimal between 0 and 1.0 with 1.0 as full completion of the route. The rest infraction components, which described the occurrences of collisions, running red lights, and lane deviations (i.e., $N_{c}$, $N_{l}$ and $N_{s}$ respectively) per run, were all nonnegative numbers. We used a right-tailed $t$-test to test against the alternative hypothesis that the population mean of the simulated driving score for BCI-JJ was greater than that for the motor intact control group of 20 participants. We used a both-tailed $t$-test to test against the alternative hypothesis that the population means for each infraction component were not equal between BCI-JJ and the motor intact control group. If $p<0.05$, the metrics had a significant impact on the performance.
@@
Table S12: Route lengths for three driving tasks. The route lengths were estimated from route maps for the Mcity teledriving task, the obstacle-course teledriving task, and the simulated town driving task. CW: clockwise. CCW: counterclockwise.
Task Name Route Option Route Length (mi)
Mcity Teledriving Mcity Route 1 0.142
Mcity Route 2 0.440
Mcity Route 3 0.197
Mcity Route 4 0.165
Obstacle-course Teledriving A course loop with a lane switch (either CW or CCW) 0.116
A course loop without any lane switch (either CW or CCW) 0.115
Simulated Town Driving A virtual town route 0.967
Table S13: Safety criteria for driving tasks. We set up the safety criteria for all three driving tasks to include some basic components of a standard driving test, in accordance with part of the evaluation and metrics from CARLA Autonomous Driving Leaderboard 2.0 and part of the checklist in the “Driving Performance Evaluation Score Sheet” from the California Department of Motor Vehicles (DMV) [44].
Infraction Infraction Corresponding California DMV Related
Components Descriptions Safety Criteria Performance Checklist
Collisions Collisions with other vehicles (simulation), or static elements (e.g., cones, bumps) No collisions with any vehicle or static elements (e.g., bumps, buildings) in the simulator. No actual or potential collisions with any manual vehicle reverse during teledriving. • “Keep an adequate space cushion between vehicles” • “Yield to oncoming traffic when appropriate”
Lane Deviations Off-lane driving from a straight or curved lane (e.g., a turn, a roundabout) In-lane driving on a straight or curved lane (e.g., a turn, a roundabout) for all four wheels. • “Begin and end turns in correct lane” • “Do not cut turns too short” • “Do not make turns too wide” • “Keep in center of lane” • “Do not under-steer” • “Do not over-steer”
Running Red Traffic Lights or Stop Signs Running a red light or a stop sign without a full stop behind the limit line An in-time full stop right behind the limit line at a stop sign or while waiting at the red traffic light. • “Make full stops behind limit lines” • “Do not make unnecessary stops” • “Make smooth safe stops”
@@
Caption for Movie S1: BCI teledriving of a commercially available Ford Mustang Mach-E via Route 1 in Mcity by BCI-JJ using the single-effector cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle (bottom-left) would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 1 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S2: BCI teledriving of a commercially available Ford Mustang Mach-E via Route 2 in Mcity by BCI-JJ using the single-effector cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle (bottom-left) would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 2 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S3: BCI teledriving of a commercially available Ford Mustang Mach-E via Route 3 in Mcity by BCI-JJ using the single-effector cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 3 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S4: BCI teledriving of a commercially available Ford Mustang Mach-E via Route 4 in Mcity by BCI-JJ using the single-effector cursor control. We used a camera mounted outside the front windshield of the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. The cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle. If the vehicle was about to collide with some dangerous obstacle and/or experienced more than two seconds of aggregated lags, the safety driver inside the vehicle (bottom-left) would take over the vehicle control for a few seconds to move it back to the right track before the BCI teledriving could resume. Route 4 map (bottom-right) has the current segment labeled in dark purple. The video is played back at 4 times the speed of the original recording.
@@
Caption for Movie S5: BCI teledriving of a commercially available Ford Mustang Mach-E through the obstacle course clockwise with a lane-switch segment by BCI-JJ using the single-effector cursor control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at twice the speed of the original recording.
@@
Caption for Movie S6: BCI teledriving of a commercially available Ford Mustang Mach-E through the obstacle course clockwise without any lane switch by BCI-JJ using the single-effector cursor control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at twice the speed of the original recording.
@@
Caption for Movie S7: BCI teledriving of a commercially available Ford Mustang Mach-E through the obstacle course counterclockwise with a lane-switch segment by BCI-JJ using the single-effector cursor control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at twice the speed of the original recording.
@@
Caption for Movie S8: BCI teledriving of a commercially available Ford Mustang Mach-E through the obstacle course counterclockwise without any lane switch by BCI-JJ using the single-effector cursor control. The example run had its teledriving score closest to the average scores for this route option. We used the camera mounted within the vehicle in Michigan for continuous recording of the view ahead and real-time video feedback to BCI-JJ (top-right) in California. As shown on the overlay (bottom-middle), the cursor movement was decoded with BCI-JJ’s right thumb for steering and speed control of the vehicle, whereas the click information was decoded with his left index finger for its full-stop braking. The route map (bottom-right) has the chosen direction of the loop and the chosen lane switch option labeled with the black arrow. The video is played back at twice the speed of the original recording.
Caption for Movie S9: Side-by-side comparisons of the average simulated town driving performance between BCI-JJ and a motor intact participant using the bimanual cursor-and-click control. We showed two example runs from BCI-JJ and one motor intact participant on the town route modified from the CARLA Leaderboard 2.0. Among all the collected runs, these two example runs had their simulated driving scores closest to the average scores for BCI-JJ and the motor intact control group. We retrieved the cursor movement and click information from either the BCI decoder for BCI-JJ or the joystick for motor intact participants. With the overlay instance floating above the vehicle simulation window, each participant got real-time visual feedback of their cursor and click control, the view ahead of the vehicle, and the side views. The cursor movement with the right thumb controlled steering and speed of the virtual vehicle, whereas the clicks with the left index finger controlled its full-stop braking. Each route map (top) corresponding to each participant’s driving recording has the current segment labeled in dark purple. The video is played back at 16 times the speed of the original recording.